OpenStudy (anonymous):

-48x^2+3x^4 find the zeros of the polynomial

OpenStudy (anonymous):

you equate this equation to zero -48x^2+3x^4=0 3x^2(-16+x^2)=0 then from here either 3x^2=0 or x^2-16=0 then x1=x2=0 and x^2=16 hence x=+4,+4 then the zeros of the equation are 0,0,-4,+4

OpenStudy (anonymous):

iam sorry, x^2=16 hence x=-4,+4 not +4,+4